OpenStudy (anonymous):

just to get it confirm -30=-48 ????? is it -78?

OpenStudy (anonymous):

-30=-48? That is impossible.

OpenStudy (anonymous):

What are you asking? -30 - -48? If so, that would be 18..

OpenStudy (anonymous):

Solve -6(5-2z) + 5z = 8z - 3(16-3z).Please show all your work?

OpenStudy (anonymous):

-6(5-2z) + 5z = 8z - 3(16-3z) -30 + 12z + 5z = 8z -48 +6z 17z= - 18 + 14z 3z = -18 z= -6